OpenStudy (anonymous):

George bought a new graduated cylinder for his chemistry class. It holds 728.97 milliliters of liquid. If the cylinder has a radius of 4.7 cm, then how tall is the cylinder? (1 milliliter is equal to 1 cubic centimeter) 155.1 cm 24,056.01 cm 77.55 cm 33 cm i need help asap

OpenStudy (calculusxy):

The formula for finding the volume of a cylinder is: \[V = Base \times height\] \[Base\] represents the area of the circle (which is the base): \[\pi \times r^2\] So here you have: \[(\pi \times r^2) \times height\]

OpenStudy (anonymous):

33?

OpenStudy (calculusxy):

Wait, I am just going to calculate it.

OpenStudy (calculusxy):

How did you get 33?

OpenStudy (anonymous):

h=v/pie*r^2

OpenStudy (anonymous):

it told me not to use 3.12 as pi

OpenStudy (calculusxy):

Pi is 3.14 not 3.12. So what number are you using for pi, or are you using the symbol on the calculator?

OpenStudy (anonymous):

thats what i meant. sorry. i was using the symbol on the calculator. the question said 33 was right.

OpenStudy (calculusxy):

Okay great!!! As long as you are correct :)

OpenStudy (anonymous):

thanks! :)